OpenStudy (anonymous):

How would solve this 5x^2 = –45x

OpenStudy (callisto):

5x^2 = –45x 5x^2 +45x = 0 Take out common factors 5x(x+9) =0 So 5x =0 or x+9 = 0 x =0 or x=-9 Do you understand that?

OpenStudy (anonymous):

hold on how did you get four answers

OpenStudy (callisto):

nah.. i've got 2 only :S

OpenStudy (callisto):

5x =0 or x+9 = 0 <- solve it and you get the next step x =0 or x=-9

OpenStudy (anonymous):

ohhhhhhh ok, i get it, thanks